# Tiny Recursive Models on ARC-AGI-1: Inductive Biases, Identity Conditioning, and Test-Time Compute

### v1

@@ 1 Introduction
The Abstraction and Reasoning Corpus (ARC) was introduced as a benchmark for evaluating generalizable, human-like reasoning in artificial systems [chollet2019measure]. ARC and its successors, ARC-AGI-1 and ARC-AGI-2, are explicitly designed to be difficult for standard large-scale pattern-matching methods, even when such methods succeed on more conventional vision and language tasks. Each ARC task consists of a small set of input–output grid pairs and requires the learner to infer and apply a transformation rule to new inputs.
Large language models (LLMs) have recently demonstrated non-trivial performance on ARC-style benchmarks when equipped with task-specific tokenization, chain-of-thought prompting, and test-time training or adaptation [wei2022chain, snell2024scaling]. However, these systems typically have very high parameter counts and incur significant inference latency and memory costs, especially when they rely on multiple sampled trajectories or search-based decoding.
Tiny Recursive Models (TRM) were proposed as a contrasting approach that combines a small parameter footprint with recursive inference over latent states [jolicoeurmartineau2025trm]. In the original study, a 7M-parameter TRM trained with heavy data augmentation and deep supervision reportedly achieves strong performance on ARC-AGI-1 and ARC-AGI-2 while operating at much lower memory and latency than typical LLMs. This has led to interest in TRM as a “tiny but capable” reasoning architecture.
@@ 2.1 ARC and ARC-AGI Benchmarks
ARC was proposed as a benchmark for general intelligence that emphasizes few-shot abstraction and compositional reasoning rather than large-scale supervised learning [chollet2019measure]. ARC-AGI-1 and ARC-AGI-2 extend this setting and have been adopted in challenge settings and leaderboards, including recent ARC Prize competitions and public leaderboards [arcprize2024arcagi1]. Public reports and leaderboards show that a range of systems, from program synthesis pipelines to frontier LLMs with specialized prompting and test-time training, achieve non-trivial but still subhuman accuracy, with a notable drop in performance when moving from ARC-AGI-1 to ARC-AGI-2 [chollet2025arcagi2].
@@ 2.2 Recursive Reasoning Models and Tiny Recursive Models
Hierarchical Reasoning Models (HRM) were proposed as a class of small recurrent architectures operating over latent states to solve structured puzzles such as Sudoku, mazes, and ARC-style tasks [wang2025hrm]. HRM combines multiple recurrent modules operating at different “frequencies” with deep supervision across many steps and relies heavily on data augmentation. Follow-up analyses have suggested that deep supervision and augmentation contribute substantially to performance, and that the theoretical justification for some components may not strictly hold in practice [arcprize2025hrmanalysis].
Tiny Recursive Models (TRM) build directly on this line of work, simplifying the architecture while improving performance on several benchmarks [jolicoeurmartineau2025trm]. TRM replaces the dual-network hierarchy with a single small network, makes the recursion mechanism more explicit, and incorporates architectural choices that favor efficiency on small grids. The original TRM paper reports strong performance on Sudoku Extreme, Maze Hard, ARC-AGI-1, and ARC-AGI-2, and emphasizes that a 7M-parameter model can outperform many larger systems when combined with heavy augmentation, deep supervision, and test-time ensembling.
Subsequent technical reports have examined TRM from complementary angles, including test-time adaptation and fine-tuning on new tasks [mcgovern2025trelis]. These works generally treat TRM as a black box and focus on end-to-end performance. In contrast, our goal is to analyze the behavior of one fixed verification checkpoint and to isolate the roles of puzzle identity, recursion depth, and augmentation in that specific instance.
@@ 2.3 Large Language Models, Chain-of-Thought, and Test-Time Compute
Large language models have shown substantial progress on reasoning benchmarks, including ARC-like tasks, when equipped with chain-of-thought prompting and test-time adaptation. Chain-of-thought methods explicitly elicit intermediate reasoning steps, often improving performance at the cost of increased inference time [wei2022chain]. More recent work formalizes the notion of test-time compute and shows that allocating additional sampling, search, or optimization budget at inference time can sometimes be as impactful as scaling model parameters [snell2024scaling].
@@ 4.1 Motivation
The TRM verification pipeline uses 1000 augmented variations of each test puzzle and selects the final answer via majority voting. This procedure can substantially improve performance even for models that are individually weak, similar to how self-consistency and ensemble sampling can improve LLM accuracy [wei2022chain, snell2024scaling]. It is therefore important to separate the contribution of the base model from the contribution of test-time compute.
@@ 8.2 Design
We train a Llama 3 8B Instruct model [grattafiori2024llama3] using QLoRA [dettmers2023qlora] on the canonical ARC-AGI-1 training set. The setup is deliberately simple:
@@ 10 Limitations
    Restricted datasets and tasks. We focus exclusively on ARC-AGI-1 for both evaluation and most training experiments. We do not analyze ARC-AGI-2, Sudoku Extreme, Maze Hard, or other benchmarks where TRM has been reported to perform well [jolicoeurmartineau2025trm]. Consequently, we cannot directly explain why the full TRM training pipeline achieves higher relative gains on some tasks or why performance differs between ARC-AGI-1 and ARC-AGI-2. Any connections to those results are conjectural.

### v2

@@ 1 Introduction
The Abstraction and Reasoning Corpus (ARC) was introduced as a benchmark for evaluating generalizable, human-like reasoning in artificial systems [4]. ARC and its successors, ARC-AGI-1 and ARC-AGI-2, are explicitly designed to be difficult for standard large-scale pattern-matching methods, even when such methods succeed on more conventional vision and language tasks. Each ARC task consists of a small set of input–output grid pairs and requires the learner to infer and apply a transformation rule to new inputs.
Large language models (LLMs) have recently demonstrated non-trivial performance on ARC-style benchmarks when equipped with task-specific tokenization, chain-of-thought prompting, and test-time training or adaptation [11, 9]. However, these systems typically have very high parameter counts and incur significant inference latency and memory costs, especially when they rely on multiple sampled trajectories or search-based decoding.
Tiny Recursive Models (TRM) were proposed as a contrasting approach that combines a small parameter footprint with recursive inference over latent states [7]. In the original study, a 7M-parameter TRM trained with heavy data augmentation and deep supervision reportedly achieves strong performance on ARC-AGI-1 and ARC-AGI-2 while operating at much lower memory and latency than typical LLMs. This has led to interest in TRM as a “tiny but capable” reasoning architecture.
@@ 2.1 ARC and ARC-AGI Benchmarks
ARC was proposed as a benchmark for general intelligence that emphasizes few-shot abstraction and compositional reasoning rather than large-scale supervised learning [4]. ARC-AGI-1 and ARC-AGI-2 extend this setting and have been adopted in challenge settings and leaderboards, including recent ARC Prize competitions and public leaderboards [1]. Public reports and leaderboards show that a range of systems, from program synthesis pipelines to frontier LLMs with specialized prompting and test-time training, achieve non-trivial but still subhuman accuracy, with a notable drop in performance when moving from ARC-AGI-1 to ARC-AGI-2 [3].
@@ 2.2 Recursive Reasoning Models and Tiny Recursive Models
Hierarchical Reasoning Models (HRM) were proposed as a class of small recurrent architectures operating over latent states to solve structured puzzles such as Sudoku, mazes, and ARC-style tasks [10]. HRM combines multiple recurrent modules operating at different “frequencies” with deep supervision across many steps and relies heavily on data augmentation. Follow-up analyses have suggested that deep supervision and augmentation contribute substantially to performance, and that the theoretical justification for some components may not strictly hold in practice [2].
Tiny Recursive Models (TRM) build directly on this line of work, simplifying the architecture while improving performance on several benchmarks [7]. TRM replaces the dual-network hierarchy with a single small network, makes the recursion mechanism more explicit, and incorporates architectural choices that favor efficiency on small grids. The original TRM paper reports strong performance on Sudoku Extreme, Maze Hard, ARC-AGI-1, and ARC-AGI-2, and emphasizes that a 7M-parameter model can outperform many larger systems when combined with heavy augmentation, deep supervision, and test-time ensembling.
Subsequent technical reports have examined TRM from complementary angles, including test-time adaptation and fine-tuning on new tasks [8]. These works generally treat TRM as a black box and focus on end-to-end performance. In contrast, our goal is to analyze the behavior of one fixed verification checkpoint and to isolate the roles of puzzle identity, recursion depth, and augmentation in that specific instance.
@@ 2.3 Large Language Models, Chain-of-Thought, and Test-Time Compute
Large language models have shown substantial progress on reasoning benchmarks, including ARC-like tasks, when equipped with chain-of-thought prompting and test-time adaptation. Chain-of-thought methods explicitly elicit intermediate reasoning steps, often improving performance at the cost of increased inference time [11]. More recent work formalizes the notion of test-time compute and shows that allocating additional sampling, search, or optimization budget at inference time can sometimes be as impactful as scaling model parameters [9].
@@ 4.1 Motivation
The TRM verification pipeline uses 1000 augmented variations of each test puzzle and selects the final answer via majority voting. This procedure can substantially improve performance even for models that are individually weak, similar to how self-consistency and ensemble sampling can improve LLM accuracy [11, 9]. It is therefore important to separate the contribution of the base model from the contribution of test-time compute.
@@ 8.2 Design
We train a Llama 3 8B Instruct model [6] using QLoRA [5] on the canonical ARC-AGI-1 training set. The setup is deliberately simple:
@@ 10 Limitations
    Restricted datasets and tasks. We focus exclusively on ARC-AGI-1 for both evaluation and most training experiments. We do not analyze ARC-AGI-2, Sudoku Extreme, Maze Hard, or other benchmarks where TRM has been reported to perform well [7]. Consequently, we cannot directly explain why the full TRM training pipeline achieves higher relative gains on some tasks or why performance differs between ARC-AGI-1 and ARC-AGI-2. Any connections to those results are conjectural.